# Evolutionarily Stable Stackelberg Equilibrium

### 19 Mar 2026 v1

@@ I Introduction
Now we present the Stackelberg evolutionary game setting. The leader’s set of pure strategies is $\mathcal{M}=\{1,\ldots,m\}$ and the follower phenotypes are $\mathcal{P}=\{1,\ldots,n\}.$ The leader’s utility function is given by $\mathbf{A_{L}}\in\mathbb{R}^{m\times n}$, where $\mathbf{A_{L}}(\ell,i)$ gives the payoff to the leader when the leader plays pure strategy $\ell\in\mathcal{M}$ and the follower population state is phenotype $i\in\mathcal{P}.$ The leader’s expected utility when the leader plays mixed strategy $\boldsymbol{\sigma}$ and the follower plays mixed strategy $\mathbf{x}$ is $U_{L}(\boldsymbol{\sigma},\mathbf{x})=\sum_{\ell=1}^{m}\sum_{i=1}^{n}\sigma_{\ell}x_{i}\mathbf{A_{L}}(\ell,i).$ The follower payoff tensor is $\mathbf{A_{F}}\in\mathbb{R}^{m\times n\times n},$ where $\mathbf{A_{F}}(\ell,i,j)$ is the payoff (fitness) to a follower of phenotype $i\in\mathcal{P}$ when the leader plays pure strategy $\ell\in\mathcal{M}$ and the interacting follower has phenotype $j\in\mathcal{P}.$ For each fixed $\ell\in\mathcal{M},$ $\mathbf{A_{F}}(\ell,\cdot,\cdot)$ is an $n\times n$ symmetric payoff matrix. After the leader commits to mixed strategy $\boldsymbol{\sigma}\in\Delta(\mathcal{M})$, the induced follower payoff matrix is $\mathbf{B}^{\boldsymbol{\sigma}}\in\mathbb{R}^{n\times n}$ with $\mathbf{B}^{\boldsymbol{\sigma}}=\sum_{\ell=1}^{m}\sigma_{\ell}\mathbf{A_{F}}(\ell,\cdot,\cdot).$ The follower population then plays a symmetric evolutionary game with payoff matrix $\mathbf{B}^{\boldsymbol{\sigma}}.$ A population state $\mathbf{x}\in\Delta(\mathcal{P})$ is admissible for $\boldsymbol{\sigma}$ iff it is an ESS of $\mathbf{B}^{\boldsymbol{\sigma}}.$ Let $\mathcal{E}(\boldsymbol{\sigma})$ denote the set of ESSs of $\mathbf{B}^{\boldsymbol{\sigma}},$ and let $g^{\boldsymbol{\sigma}}:\Delta(\mathcal{P})\rightarrow\mathbb{R}$ denote the follower’s ESS selection function. Finally let $\mathcal{G}(\boldsymbol{\sigma})=\operatorname*{arg\,max}_{\mathbf{x}\in\mathcal{E}(\boldsymbol{\sigma})}g^{\boldsymbol{\sigma}}(\mathbf{x}).$
@@
We now present our general definition for evolutionarily stable Stackelberg equilibrium (SESS) in Definition 4. We also define two natural special cases. In the first case, the followers play the ESS that is most beneficial to the leader. We refer to this as an optimistic evolutionarily stable Stackelberg equilibrium (OSESS). In this case we have $g^{\boldsymbol{\sigma}}(\mathbf{x})=U_{L}(\boldsymbol{\sigma},\mathbf{x}).$ We also consider pessimistic evolutionarily stable Stackelberg equilibrium (PSESS), in which the followers play the ESS that is worst for the leader. While the followers are not rationally trying to harm the leader, we can view PSESS as SESS that is robust to worst-case evolutionary stability. For PSESS we have $g^{\boldsymbol{\sigma}}(\mathbf{x})=-U_{L}(\boldsymbol{\sigma},\mathbf{x}).$ OSESS and PSESS are defined below in Definitions 56. For OSESS we have $\mathcal{G}(\boldsymbol{\sigma})=\operatorname*{arg\,max}_{\mathbf{x}\in\mathcal{E}(\boldsymbol{\sigma})}U_{L}(\boldsymbol{\sigma},\mathbf{x})$, and for PSESS $\mathcal{G}(\boldsymbol{\sigma})=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{E}(\boldsymbol{\sigma})}U_{L}(\boldsymbol{\sigma},\mathbf{x}).$ SESS refines Stackelberg equilibrium by restricting the follower response to strategies that are evolutionarily stable in the induced population game.
@@
$\boldsymbol{\sigma}^{*}\in\operatorname*{arg\,max}_{\boldsymbol{\sigma}\in\Delta(\mathcal{M})}\;\min_{\mathbf{x}\in\mathcal{E}(\boldsymbol{\sigma})}U_{L}(\boldsymbol{\sigma},\mathbf{x})\quad\text{and}\quad\mathbf{x}^{*}\in\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{E}(\boldsymbol{\sigma}^{*})}U_{L}(\boldsymbol{\sigma}^{*},\mathbf{x}).$
@@ II Continuous Stackelberg–ESS
For a fixed leader decision $\mathbf{m}$, we define the set of admissible follower responses $\mathcal{E}(\mathbf{m})\subseteq\mathcal{Z}$ as the set of eco–evolutionarily stable outcomes under $\mathbf{m}$. Informally, a follower outcome $\mathbf{z}^{*}=(\mathbf{x}^{*},\mathbf{u}^{*})$ is admissible if it satisfies any required ecological equilibrium conditions and is stable against invasion by rare mutant phenotypes in the induced eco–evolutionary system. We assume throughout that $\mathcal{E}(\mathbf{m})$ is nonempty for the leader decisions under consideration. When multiple evolutionarily stable outcomes exist for the same leader decision, we model their selection via a correspondence $\mathcal{G}(\mathbf{m})\subseteq\mathcal{E}(\mathbf{m})$.
We now define continuous analogues of Stackelberg–ESS.
@@
$\mathbf{m}^{*}\in\operatorname*{arg\,max}_{\mathbf{m}\in\mathcal{M}}\ \min_{\mathbf{z}\in\mathcal{E}(\mathbf{m})}Q(\mathbf{m},\mathbf{z})\quad\text{and}\quad\mathbf{z}^{*}\in\operatorname*{arg\,min}_{\mathbf{z}\in\mathcal{E}(\mathbf{m}^{*})}Q(\mathbf{m}^{*},\mathbf{z}).$
First we present the Stackelberg evolutionary game formulation of cancer treatment from prior work [18]. In this formulation (Equation 1), the objective Q corresponds to the quality of life function that is maximized by the leader. For each cell type $i$, there is a fitness function $G_{i}(u_{i},\mathbf{m},\mathbf{x})$ that the follower is trying to maximize. We assume that the dynamics of the population $\mathbf{x}$ are governed by $\dot{\mathbf{x}}=\mathbf{G}(t)\mathbf{x}.$ In order to ensure that we are in equilibrium of the ecological dynamics we must have that $\dot{x_{i}}=0$ for all $i.$ Note that in this formulation the follower is playing a best response by maximizing the fitness function. By contrast, in Stackelberg-ESS we assume that the behavior of the cancer cells is determined by evolutionary stability conditions as opposed to explicit utility maximization. An algorithm for solving this problem has been presented that is based on a nonconvex quadratic program formulation [8]. This algorithm has been demonstrated to quickly compute a global optimal solution on a realistic example problem proposed by prior work [18].
@@ III Algorithms
$\max_{u}$ $G_{i}(u,\mathbf{m}^{\star},\mathbf{x}^{\star})$
s.t. $0\leq u\leq 1$
@@ V Conclusion
We define SESS both for discrete normal-form games, as well continuous-trait games, which arise naturally in biological and control applications such as cancer treatment. In our general formulation, the followers play any ESS in the subgame induced by the leader’s strategy. We also consider the special cases where the followers play the ESS that is most beneficial to the leader (OSESS) and the ESS that is worst for the leader (PSESS). OSESS is perhaps the easiest case of SESS to compute since in a sense all players are aligned in maximizing the leader’s payoff (subject to the followers playing an ESS), which simplifies the optimization to a single outer maximization problem. We present algorithms for computing OSESS in both the discrete and continuous settings. If we view the follower strategies as population frequencies, then our model is applicable to settings with a large number of follower players; e.g., there may be a large number of cancer cells where each cell has one of a small number of possible phenotypes.

### 25 Mar 2026 v2

@@ I Introduction
Now we present the Stackelberg evolutionary game setting. The leader’s set of pure strategies is $\mathcal{M}=\{1,\ldots,m\}$ and the follower phenotypes are $\mathcal{P}=\{1,\ldots,n\}.$ The leader’s utility function is given by $\mathbf{A_{L}}\in\mathbb{R}^{m\times n}$, where $\mathbf{A_{L}}(\ell,i)$ gives the payoff to the leader when the leader plays pure strategy $\ell\in\mathcal{M}$ and the follower population state is phenotype $i\in\mathcal{P}.$ The leader’s expected utility when the leader plays mixed strategy $\boldsymbol{\sigma}$ and the follower plays mixed strategy $\mathbf{x}$ is $U_{L}(\boldsymbol{\sigma},\mathbf{x})=\sum_{\ell=1}^{m}\sum_{i=1}^{n}\sigma_{\ell}x_{i}\mathbf{A_{L}}(\ell,i).$ The follower payoff tensor is $\mathbf{A_{F}}\in\mathbb{R}^{m\times n\times n},$ where $\mathbf{A_{F}}(\ell,i,j)$ is the payoff (fitness) to a follower of phenotype $i\in\mathcal{P}$ when the leader plays pure strategy $\ell\in\mathcal{M}$ and the interacting follower has phenotype $j\in\mathcal{P}.$ For each fixed $\ell\in\mathcal{M},$ $\mathbf{A_{F}}(\ell,\cdot,\cdot)$ is an $n\times n$ symmetric payoff matrix. After the leader commits to mixed strategy $\boldsymbol{\sigma}\in\Delta(\mathcal{M})$, the induced follower payoff matrix is $\mathbf{B}^{\boldsymbol{\sigma}}\in\mathbb{R}^{n\times n}$ with $\mathbf{B}^{\boldsymbol{\sigma}}=\sum_{\ell=1}^{m}\sigma_{\ell}\mathbf{A_{F}}(\ell,\cdot,\cdot).$ The follower population then plays a symmetric evolutionary game with payoff matrix $\mathbf{B}^{\boldsymbol{\sigma}}.$ A population state $\mathbf{x}\in\Delta(\mathcal{P})$ is admissible for $\boldsymbol{\sigma}$ iff it is an ESS of $\mathbf{B}^{\boldsymbol{\sigma}}.$ Let $\mathcal{E}(\boldsymbol{\sigma})$ denote the set of ESSs of $\mathbf{B}^{\boldsymbol{\sigma}},$ and let $g^{\boldsymbol{\sigma}}:\Delta(\mathcal{P})\rightarrow\mathbb{R}$ denote the follower’s ESS selection function. Finally let $\mathcal{G}(\boldsymbol{\sigma})=\operatorname*{arg\,max}_{\mathbf{x}\in\mathcal{E}(\boldsymbol{\sigma})}g^{\boldsymbol{\sigma}}(\mathbf{x}).$ The leader maximizes utility anticipating that the follower outcome is selected from the correspondence $\mathcal{G}(\boldsymbol{\sigma}).$
@@
We now present our general definition for evolutionarily stable Stackelberg equilibrium (SESS) in Definition 4. We also define two natural special cases. In the first case, the followers play the ESS that is most beneficial to the leader. We refer to this as an optimistic evolutionarily stable Stackelberg equilibrium (OSESS). In this case we have $g^{\boldsymbol{\sigma}}(\mathbf{x})=U_{L}(\boldsymbol{\sigma},\mathbf{x}).$ We also consider pessimistic evolutionarily stable Stackelberg equilibrium (PSESS), in which the followers play the ESS that is worst for the leader. While the followers are not rationally trying to harm the leader, we can view PSESS as SESS that is robust to worst-case evolutionary stability. For PSESS we have $g^{\boldsymbol{\sigma}}(\mathbf{x})=-U_{L}(\boldsymbol{\sigma},\mathbf{x}).$ OSESS and PSESS are defined below in Definitions 56. For OSESS we have $\mathcal{G}(\boldsymbol{\sigma})=\operatorname*{arg\,max}_{\mathbf{x}\in\mathcal{E}(\boldsymbol{\sigma})}U_{L}(\boldsymbol{\sigma},\mathbf{x})$, and for PSESS $\mathcal{G}(\boldsymbol{\sigma})=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{E}(\boldsymbol{\sigma})}U_{L}(\boldsymbol{\sigma},\mathbf{x}).$ It is clear that OSESS and PSESS are refinements of SESS (since they specify a particular function $g^{\boldsymbol{\sigma}}(\mathbf{x})$). SESS can be viewed as a refinement of Stackelberg equilibrium in which the follower response correspondence is restricted to evolutionarily stable strategies in the induced population game (Proposition 1).
@@
$\boldsymbol{\sigma}^{*}\in\operatorname*{arg\,max}_{\boldsymbol{\sigma}\in\Delta(\mathcal{M})}\;\min_{\mathbf{x}\in\mathcal{E}(\boldsymbol{\sigma})}U_{L}(\boldsymbol{\sigma},\mathbf{x})\text{ and }\mathbf{x}^{*}\in\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{E}(\boldsymbol{\sigma}^{*})}U_{L}(\boldsymbol{\sigma}^{*},\mathbf{x}).$
Proposition 1
Every evolutionarily stable Stackelberg equilibrium $(\boldsymbol{\sigma}^{*},\boldsymbol{x}^{*})$ is a Stackelberg equilibrium.
Proof:
Suppose that $(\boldsymbol{\sigma}^{*},\boldsymbol{x}^{*})$ is an evolutionarily stable Stackelberg equilibrium. By Definition 4,
$\boldsymbol{\sigma}^{*}\in\arg\max_{\boldsymbol{\sigma}\in\Delta(M)}\max_{\boldsymbol{x}\in\mathcal{G}(\boldsymbol{\sigma})}U_{L}(\boldsymbol{\sigma},\boldsymbol{x})\quad\text{and}\quad\boldsymbol{x}^{*}\in\mathcal{G}(\boldsymbol{\sigma}^{*}).$
Since $\mathcal{G}(\boldsymbol{\sigma}^{*})\subseteq\mathcal{E}(\boldsymbol{\sigma}^{*})$, it follows that
$\boldsymbol{x}^{*}\in\mathcal{E}(\boldsymbol{\sigma}^{*}),$
so the follower outcome $\boldsymbol{x}^{*}$ is an evolutionarily stable strategy in the induced follower game under $\boldsymbol{\sigma}^{*}$.
Moreover, $\boldsymbol{\sigma}^{*}$ maximizes the leader’s utility over all strategies, anticipating that the follower response is selected from $\mathcal{G}(\boldsymbol{\sigma})$. Since $\boldsymbol{x}^{*}\in\mathcal{G}(\boldsymbol{\sigma}^{*})$, the pair $(\boldsymbol{\sigma}^{*},\boldsymbol{x}^{*})$ satisfies the Stackelberg optimality and consistency conditions. Hence it is a Stackelberg equilibrium. ∎
@@ II Continuous Stackelberg–ESS
For a fixed leader decision $\mathbf{m}$, we define the set of admissible follower responses $\mathcal{E}(\mathbf{m})\subseteq\mathcal{Z}$ as the set of eco–evolutionarily stable outcomes under $\mathbf{m}$. Informally, a follower outcome $\mathbf{z}^{*}=(\mathbf{x}^{*},\mathbf{u}^{*})$ is admissible if it satisfies any required ecological equilibrium conditions and is stable against invasion by rare mutant phenotypes in the induced eco–evolutionary system. When multiple evolutionarily stable outcomes exist for the same leader decision, we model their selection via a correspondence $\mathcal{G}(\mathbf{m})\subseteq\mathcal{E}(\mathbf{m})$. We assume that $\mathcal{E}(\boldsymbol{m})\neq\emptyset$ for all $\boldsymbol{m}\in M$, so that the follower response correspondence is well-defined.
We now define continuous analogues of Stackelberg–ESS. It is again clear that in the continuous setting OSESS and PSESS are refinements of SESS, and SESS is a refinement of Stackelberg equilibrium in which follower responses are restricted to evolutionarily stable strategies.
@@
$\mathbf{m}^{*}\in\operatorname*{arg\,max}_{\mathbf{m}\in\mathcal{M}}\ \min_{\mathbf{z}\in\mathcal{E}(\mathbf{m})}Q(\mathbf{m},\mathbf{z})\text{ and }\mathbf{z}^{*}\in\operatorname*{arg\,min}_{\mathbf{z}\in\mathcal{E}(\mathbf{m}^{*})}Q(\mathbf{m}^{*},\mathbf{z}).$
Proposition 2
Every continuous evolutionarily stable Stackelberg equilibrium is a Stackelberg equilibrium.
Proof:
Suppose that $(\boldsymbol{m}^{*},\boldsymbol{z}^{*})$ is a continuous evolutionarily stable Stackelberg equilibrium. By Definition 7,
$\boldsymbol{m}^{*}\in\arg\max_{\boldsymbol{m}\in M}\max_{\boldsymbol{z}\in\mathcal{G}(\boldsymbol{m})}Q(\boldsymbol{m},\boldsymbol{z})\quad\text{and}\quad\boldsymbol{z}^{*}\in\mathcal{G}(\boldsymbol{m}^{*}).$
Since $\mathcal{G}(\boldsymbol{m}^{*})\subseteq\mathcal{E}(\boldsymbol{m}^{*})$, it follows that $\boldsymbol{z}^{*}\in\mathcal{E}(\boldsymbol{m}^{*}),$ so the follower outcome $\boldsymbol{z}^{*}$ is eco–evolutionarily stable under $\boldsymbol{m}^{*}$. Moreover, $\boldsymbol{m}^{*}$ maximizes the leader’s objective over all decisions, anticipating that the follower response is selected from $\mathcal{G}(\boldsymbol{m})$. Since $\boldsymbol{z}^{*}\in\mathcal{G}(\boldsymbol{m}^{*})$, the pair $(\boldsymbol{m}^{*},\boldsymbol{z}^{*})$ satisfies the Stackelberg optimality and consistency conditions. Hence it is a Stackelberg equilibrium. ∎
We now present the Stackelberg evolutionary game formulation of cancer treatment from prior work [18]. In this formulation (Equation 1), the objective Q corresponds to the quality of life function that is maximized by the leader. For each cell type $i$, there is a fitness function $G_{i}(u_{i},\mathbf{m},\mathbf{x})$ that the follower is trying to maximize. We assume that the dynamics of the population $\mathbf{x}$ are governed by $\dot{\mathbf{x}}=\mathbf{G}(t)\mathbf{x}.$ In order to ensure that we are in equilibrium of the ecological dynamics we must have that $\dot{x_{i}}=0$ for all $i.$ Note that in this formulation the follower is playing a best response by maximizing its fitness function. By contrast, in Stackelberg–ESS we assume that the behavior of the cancer cells is determined by evolutionary stability conditions rather than explicit optimization. An algorithm for solving this problem has been presented that is based on a nonconvex quadratic program formulation [8]. This algorithm has been demonstrated to quickly compute a global optimal solution on a realistic example problem proposed by prior work [18].
@@ III Algorithms
$\max_{u_{i}}$ $G_{i}(u_{i},\mathbf{m}^{\star},\mathbf{x}^{\star})$
s.t. $0\leq u_{i}\leq 1$
@@ V Conclusion
We define SESS both for discrete normal-form games, as well continuous-trait games, which arise naturally in biological and control applications such as cancer treatment. In our general formulation, the followers play an ESS in the subgame induced by the leader’s strategy that maximizes an arbitrary objective function. We also consider the special cases where the followers play the ESS that is most beneficial to the leader (OSESS) and the ESS that is worst for the leader (PSESS). OSESS is perhaps the easiest case of SESS to compute since in a sense all players are aligned in maximizing the leader’s payoff (subject to the followers playing an ESS), which simplifies the optimization to a single outer maximization problem. We present algorithms for computing OSESS in both the discrete and continuous settings. If we view the follower strategies as population frequencies, then our model is applicable to settings with a large number of follower players; e.g., there may be a large number of cancer cells where each cell has one of a small number of possible phenotypes.